# Feed Two Birds with One Scone: Exploiting Wild Data for Both Out-of-Distribution Generalization and Detection

### v1

@@ 1 Introduction
These distributional shift scenarios give rise to the importance of two problems: OOD generalization, which focuses on the covariate shift problem (Gulrajani & Lopez-Paz, 2020; Koh et al., 2021; Ye et al., 2022), and OOD detection, which targets semantic shift (Hendrycks & Gimpel, 2017; Liu et al., 2020; Yang et al., 2021). Both problems have received increasing research attention lately, albeit have been pursued independently; as a result, existing methods are highly specialized in one task, but not capable of handling both simultaneously. This has largely impeded the wider adoption of OOD algorithms in real-world environments, which often present heterogeneous data shifts. A critical yet underexplored question thus arises:
@@
where $\mathbb{P}_{\text{in}}$, $\mathbb{P}_{\text{out}}^{\text{covariate}}$ and $\mathbb{P}_{\text{out}}^{\text{semantic}}$ denote the marginal distributions of ID, covariate-shifted OOD and semantic-shifted OOD data respectively. Such wild data is available in abundance, does not require any human annotation, and importantly, contains the true test time OOD distributions under both covariate and semantic shifts. Thus, our distributional model above offers strong generality and practicality, compared to previous works that primarily consider the semantic shift in the wild data (Katz-Samuels et al., 2022). Despite the promise, learning from such heterogeneous data is technically challenging due to the lack of clear membership (ID, Covariate-OOD, Semantic-OOD) for samples drawn from the wild data distribution $\mathbb{P}_{\text{wild}}$.
@@
Extensive experiments confirm that Scone can effectively improve both OOD generalization and detection performance. Compared to the most related baseline WOODS (Katz-Samuels et al., 2022), our method can substantially improve the OOD classification accuracy from $52.76\%$ to ${84.69}\%$ on covariate shifted CIFAR-10 data—a direct $\textbf{31.93}\%$ improvement. Our key contributions are:
@@ 2 Problem Setup
WOODS (Katz-Samuels et al., 2022) 88.10% 97.88% 0.194% 99.88%
@@
Table 1: WOODS (Katz-Samuels et al., 2022) displays limiting OOD generalization performance. For experiments, we use 25,000 samples from the MNIST dataset as ID, and a wild mixture dataset consisting of FashionMNIST as semantic-OOD data ($\pi_{s}$ = 0.4) and MNIST-C (Mu & Gilmer, 2019) — a covariate shifted version of MNIST — as covariate-OOD data ($\pi_{c}=0.3$).
@@
[…]
@@ 3.1 Preliminaries
Katz-Samuels et al. (2022) proposed WOODS to tackle the OOD detection problem via unlabeled wild data, which consists of the ID and semantically shifted OOD data $\mathbb{P}_{\text{wild}}:=(1-\pi)\mathbb{P}_{\text{in}}+\pi\mathbb{P}_{\text{out}}^{\text{semantic}}$. The crucial difference between our work and WOODS is whether the wild mixture data contains covariate-shifted data, which introduces new challenges not considered in prior work. As we show in this work, our formulation uniquely enables both OOD generalization and OOD detection, in one coherent framework.
@@
In particular, the OOD detector $g_{\theta}$ is defined based on the level set: $g_{\theta}(\mathbf{x})=\textsc{out}$ if ${E_{\theta}(\mathbf{x})>0}$, where the free energy $E_{\theta}(\mathbf{x}):=-\log\sum_{j=1}^{K}e^{f_{\theta}^{(j)}(\mathbf{x})}$ was shown to be an effective OOD score (Liu et al., 2020). ID data tends to have negative energy and vice versa.
@@
Limitation in OOD generalization performance of WOODS: a case study. Although WOODS can simultaneously learn an OOD detector and an ID classifier, it can perform poorly on the task of OOD generalization. To see this, we investigate the efficacy of using the constrained optimization in Equation (3) directly for our new problem setting described in Section 2. To simulate the wild data $\mathbb{P}_{\text{wild}}$, we mix a subset of MNIST training data (as $\mathbb{P}_{\text{in}}$) with MNIST-C  (Mu & Gilmer, 2019) data — a covariate-shifted version of MNIST — as $\mathbb{P}_{\text{out}}^{\text{covariate}}$, and the FashionMNIST dataset (Xiao et al., 2017) as $\mathbb{P}_{\text{out}}^{\text{semantic}}$. We train a version of WOODS on this data and summarize our findings in Table 1: interestingly, we observe a significant generalization gap between classifying ID data (97.88% in accuracy) and covariate-shifted OOD data (88.10% in accuracy). This suggests that the training objective in (Katz-Samuels et al., 2022) is indeed insufficient for the purpose of OOD generalization, despite strong OOD detection performance. This motivates our method which accounts for the task of OOD generalization on covariate-shifted data.
@@ 3.2 Proposed Method
where $\eta$ controls the margin of the OOD decision boundary w.r.t. the ID data. Note that our learning objective generalizes WOODS (Katz-Samuels et al., 2022), which corresponds to the special case of $\eta=0$ (i.e., no margin at all).
@@ 3.3 Theoretical Insights: Relationship Between OOD Generalization and Energy Margin
Suppose further that we set $\alpha=0$ and $\tau=0$, such that every ID point is classified correctly and has energy satisfying $E_{\theta}(\phi(\mathbf{x}))<\eta$. Finally, suppose that $\overline{f}_{\theta}$ is $L$-Lipschitz: this is the case for many classifier functions, such as two-layer ReLU networks with bounded variation (Parhi & Nowak, 2022). We then have the following result on the covariate-shifted points (proved in Appendix A):
@@
Under the above assumptions, if $\eta<-\log 2-\frac{1}{2}L\delta$ then each covariate-shifted point is classified correctly and is detected as semantic in.
@@ 3.4 Enforcing Margin in Practice
Solving the constrained optimization. We adopt the Augmented Lagrangian method (Hestenes, 1969) to solve our constrained optimization problem with modern neural networks. In short, the constrained optimization problem above is converted into a sequence of unconstrained optimization problems. We refer interested readers to Section 3.2 in Katz-Samuels et al. (2022) for details. We showcase the efficacy of our algorithm on the simple MNIST example in Table 1, where the proposed method improves the OOD generalization accuracy from 88.10% (WOODS) to 96.51%. Building on this encouraging result, we proceed to comprehensively evaluate our algorithm in the next section.
@@ 4.1 Experimental Setup
Following the common benchmarks in literature, we use CIFAR-10 (Krizhevsky et al., 2009) as the in-distribution data ($\mathbb{P}_{\text{in}}$). For the covariate-shifted data ($\mathbb{P}_{\text{out}}^{\text{covariate}}$), we use CIFAR-10-C (Hendrycks & Dietterich, 2018) with Gaussian additive noise for our main experiments, and provide ablations in Appendix F on other types of covariate shifts. For semantic-shifted OOD data ($\mathbb{P}_{\text{out}}^{\text{semantic}}$), we use natural image datasets: SVHN (Netzer et al., 2011), Textures (Cimpoi et al., 2014), Places365 (Zhou et al., 2017), LSUN-Crop (Yu et al., 2015), and LSUN-Resize (Yu et al., 2015). Large-scale results on the ImageNet dataset can be found in Section 4.4. Additional results on the PACS dataset (Li et al., 2017) from DomainBed is presented in Appendix E. We provide a detailed description of the datasets in Appendix C.
To simulate the wild data $\mathbb{P}_{\text{wild}}$, we mix a subset of ID data ($\mathbb{P}_{\text{in}}$) with the covariate shifted dataset ($\mathbb{P}_{\text{out}}^{\text{covariate}}$) under various $\pi_{c}\in\{0.0,0.1,0.2,0.5,0.9\}$; we default to $\pi_{c}=0.5$ for the main experiment. We keep $\pi_{s}=0.1$ to reflect the fact that we expect semantic OOD shifts ($\mathbb{P}_{\text{out}}^{\text{semantic}}$) to be encountered less frequently, which is the same value as used in (Katz-Samuels et al., 2022). We split ID datasets into two halves: we use 50% as ID training data and 50% for creating the mixture data.
@@
Training details. For CIFAR experiments and methods, we use the Wide ResNet (Zagoruyko & Komodakis, 2016) architecture with 40 layers and widen factor of 2. The model is optimized using stochastic gradient descent with Nesterov momentum (Duchi et al., 2011). We set the weight decay as 0.0005, and momentum as 0.09. We initialize the model with a pre-trained network on CIFAR-10, and then trained for 100 epochs using our method. The initial learning rate is 0.0001 and decays by a factor of 2 at epochs 50, 75, and 90. Following the previous literature (Katz-Samuels et al., 2022), we set $\alpha=0.05$, and set $\tau$ to be twice the loss of the pre-trained model. For all the experiments, we use a batch size of 128 and a dropout rate of 0.3. Our framework was implemented with PyTorch 1.8.1. All training is performed using NVIDIA GeForce RTX 2080 Ti. See Appendix B for additional experimental details, including our validation strategy for selecting $\eta$.
@@ 4.2 Results and Discussion
margin OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$
@@
Table 3: Main results: comparison with competitive OOD generalization and OOD detection methods on CIFAR-10. We run our method 3 times and report the average and std. For experiments using $\mathbb{P}_{\mathrm{wild}}$, we set $\pi_{s}=0.5$, $\pi_{c}=0.1$. For each semantic OOD dataset, we create corresponding wild mixture distribution $\mathbb{P}_{\text{wild}}:=(1-\pi_{s}-\pi_{c})\mathbb{P}_{\text{in}}+\pi_{s}\mathbb{P}_{\text{out}}^{\text{semantic}}+\pi_{c}\mathbb{P}_{\text{out}}^{\text{covariate}}$ for training and evaluating on the corresponding test dataset. ±$\pm x$ denotes the standard error, rounded to the first decimal point. Results for LSUN-R and Texture datasets are in Appendix D. (*Since all the OOD detection methods use the same model trained with the CE loss on $\mathbb{P}_{\text{in}}$, they display the same ID and OOD accuracy on CIFAR-10-C.)
@@
OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$ OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$ OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$
@@
We highlight a few observations: (1) Scone outperforms competitive post hoc OOD detection methods, including MSP (Hendrycks & Gimpel, 2017), ODIN (Liang et al., 2018), Energy (Liu et al., 2020), Mahalanobis (Lee et al., 2018),  ViM (Wang et al., 2022a), and the latest baseline KNN (Sun et al., 2022) — all of which use the same model trained with the CE loss on $\mathbb{P}_{\text{in}}$ (and hence they display the same OOD accuracy on CIFAR-10-C). As an example of our method’s improved performance, when using SVHN as $\mathbb{P}_{\text{out}}^{\text{semantic}}$ our method yields an FPR95 of 10.86% (lower is better), which outperforms the best baseline performance of 12.89%. At the same time, the OOD generalization performance is significantly improved from 75.05% (baseline CE model) to 84.69% (ours). (2) Our method also outperforms common OOD generalization baselines, including IRM (Arjovsky et al., 2019), Mixup (Zhang et al., 2018), and VREx (Krueger et al., 2021). While these methods display stronger OOD generalization performance than the ERM (or CE) baseline, they underperform ours both in terms of OOD generalization and OOD detection. [^2] Same as ours, we use the energy score in test time to perform OOD detection. (3) Lastly, we also compare to OOD detection methods utilizing $\mathbb{P}_{\text{wild}}$, including oe (Hendrycks et al., 2018), energy-regularized learning (Liu et al., 2020), and WOODS (Katz-Samuels et al., 2022), which is the latest such method to be developed. These methods are among the strongest OOD detection methods, yet display a significantly worsened OOD generalization performance. The main reason is that they make assumptions on $\mathbb{P}_{\text{wild}}$ without considering covariate OOD data.
@@
$\pi_{c}$ Method OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$
@@
In Table 4, we ablate the effect of $\pi_{c}$, which modulates the fraction of covariate OOD data in the mixture distribution $\mathbb{P}_{\text{wild}}$. For all settings, we contrast the performance without vs. with margin. The margin in each $\pi_{c}$ setting is validated using the strategy in Appendix B. Here we consistently use $\pi_{s}=0.1$, which reflects the practical scenario that the majority of test data may remain in the known classes. We primarily focus on evaluations where $\pi_{c}\neq 0$, since our problem setting uniquely introduces the covariate shift in the wild distribution. However, for completeness, we also include results when $\pi_{c}=0$. We highlight a few interesting observations: (1) without any enforced margin, the OOD generalization performance for WOODS (Katz-Samuels et al., 2022) generally degrades with increasing $\pi_{c}$. This is likely due to the fact that a larger $\pi_{c}$ translates into more severe covariate shifts. For example, when $\pi_{c}=0.9$, the OOD classification accuracy decreases to 52.84%. (2) Our method is overall more robust under large $\pi_{c}$ settings than the WOODS baseline. For instance, in a challenging case with $\pi_{c}=0.9$, Scone outperforms WOODS by 33.34%. Overall, these results demonstrate the benefits of Scone for both OOD generalization and detection.
@@
WOODS (Katz-Samuels et al., 2022) 44.46 86.49 10.50 98.22

### v2

@@ 1 Introduction
These distributional shift scenarios give rise to the importance of two problems: OOD generalization, which focuses on the covariate shift problem (gulrajani2020search; koh2021wilds; ye2022ood), and OOD detection, which targets semantic shift (hendrycks2016baseline; liu2020energy; yang2021generalized). Both problems have received increasing research attention lately, albeit have been pursued independently; as a result, existing methods are highly specialized in one task, but not capable of handling both simultaneously. This has largely impeded the wider adoption of OOD algorithms in real-world environments, which often present heterogeneous data shifts. A critical yet underexplored question thus arises:
@@
where $\mathbb{P}_{\text{in}}$, $\mathbb{P}_{\text{out}}^{\text{covariate}}$ and $\mathbb{P}_{\text{out}}^{\text{semantic}}$ denote the marginal distributions of ID, covariate-shifted OOD and semantic-shifted OOD data respectively. Such wild data is available in abundance, does not require any human annotation, and importantly, contains the true test time OOD distributions under both covariate and semantic shifts. Thus, our distributional model above offers strong generality and practicality, compared to previous works that primarily consider the semantic shift in the wild data (katz2022training). Despite the promise, learning from such heterogeneous data is technically challenging due to the lack of clear membership (ID, Covariate-OOD, Semantic-OOD) for samples drawn from the wild data distribution $\mathbb{P}_{\text{wild}}$.
@@
Extensive experiments confirm that Scone can effectively improve both OOD generalization and detection performance. Compared to the most related baseline WOODS (katz2022training), our method can substantially improve the OOD classification accuracy from $52.76\%$ to ${84.69}\%$ on covariate shifted CIFAR-10 data—a direct $\textbf{31.93}\%$ improvement. Our key contributions are:
@@ 2 Problem Setup
WOODS (katz2022training) 88.10% 97.88% 0.194% 99.88%
@@
Table 1: WOODS (katz2022training) displays limiting OOD generalization performance. For experiments, we use 25,000 samples from the MNIST dataset as ID, and a wild mixture dataset consisting of FashionMNIST as semantic-OOD data ($\pi_{s}$ = 0.4) and MNIST-C (mu2019mnist) — a covariate shifted version of MNIST — as covariate-OOD data ($\pi_{c}=0.3$).
@@
$(1)~\uparrow\text{ID-Acc}(f_{\theta}):=\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{P}_{\mathcal{X}\mathcal{Y}}}(\mathbbm{1}{\{{\widehat{y}(f_{\theta}(\mathbf{x}))}=y\}}),$
$(2)~\uparrow\text{OOD-Acc}(f_{\theta}):=\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{P}_{\text{out}}^{\text{covariate}}}(\mathbbm{1}{\{{\widehat{y}(f_{\theta}(\mathbf{x}))}=y\}}),$
@@ 3.1 Preliminaries
katz2022training proposed WOODS to tackle the OOD detection problem via unlabeled wild data, which consists of the ID and semantically shifted OOD data $\mathbb{P}_{\text{wild}}:=(1-\pi)\mathbb{P}_{\text{in}}+\pi\mathbb{P}_{\text{out}}^{\text{semantic}}$. The crucial difference between our work and WOODS is whether the wild mixture data contains covariate-shifted data, which introduces new challenges not considered in prior work. As we show in this work, our formulation uniquely enables both OOD generalization and OOD detection, in one coherent framework.
@@
In particular, the OOD detector $g_{\theta}$ is defined based on the level set: $g_{\theta}(\mathbf{x})=\textsc{out}$ if $E_{\theta}(\mathbf{x})>0$}, where the free energy $E_{\theta}(\mathbf{x}):=-\log\sum_{j=1}^{K}e^{f_{\theta}^{(j)}(\mathbf{x})}$ was shown to be an effective OOD score (liu2020energy). ID data tends to have negative energy and vice versa.
@@
Limitation in OOD generalization performance of WOODS: a case study. Although WOODS can simultaneously learn an OOD detector and an ID classifier, it can perform poorly on the task of OOD generalization. To see this, we investigate the efficacy of using the constrained optimization in Equation (3) directly for our new problem setting described in Section 2. To simulate the wild data $\mathbb{P}_{\text{wild}}$, we mix a subset of MNIST training data (as $\mathbb{P}_{\text{in}}$) with MNIST-C  (mu2019mnist) data — a covariate-shifted version of MNIST — as $\mathbb{P}_{\text{out}}^{\text{covariate}}$, and the FashionMNIST dataset (xiao2017fashion) as $\mathbb{P}_{\text{out}}^{\text{semantic}}$. We train a version of WOODS on this data and summarize our findings in Table 1: interestingly, we observe a significant generalization gap between classifying ID data (97.88% in accuracy) and covariate-shifted OOD data (88.10% in accuracy). This suggests that the training objective in (katz2022training) is indeed insufficient for the purpose of OOD generalization, despite strong OOD detection performance. This motivates our method which accounts for the task of OOD generalization on covariate-shifted data.
@@ 3.2 Proposed Method
where $\eta$ controls the margin of the OOD decision boundary w.r.t. the ID data. Note that our learning objective generalizes WOODS (katz2022training), which corresponds to the special case of $\eta=0$ (i.e., no margin at all).
@@ 3.3 Theoretical Insights: Relationship Between OOD Generalization and Energy Margin
Suppose further that we set $\alpha=0$ and $\tau=0$, such that every ID point is classified correctly and has energy satisfying $E_{\theta}(\phi(\mathbf{x}))<\eta$. Finally, suppose that $\overline{f}_{\theta}$ is $L$-Lipschitz: this is the case for many classifier functions, such as two-layer ReLU networks with bounded variation (parhi2022what). We then have the following result on the covariate-shifted points (proved in Appendix A):
@@
(Informal) Under some assumptions, if $\eta<-\log 2-\frac{1}{2}L\delta$ then each covariate-shifted point is classified correctly and is detected as semantic in.
@@ 3.4 Enforcing Margin in Practice
Solving the constrained optimization. We adopt the Augmented Lagrangian method (hestenes1969multiplier) to solve our constrained optimization problem with modern neural networks. In short, the constrained optimization problem above is converted into a sequence of unconstrained optimization problems. We refer interested readers to Section 3.2 in katz2022training for details. We showcase the efficacy of our algorithm on the simple MNIST example in Table 1, where the proposed method improves the OOD generalization accuracy from 88.10% (WOODS) to 96.51%. Building on this encouraging result, we proceed to comprehensively evaluate our algorithm in the next section.
@@ 4.1 Experimental Setup
Following the common benchmarks in literature, we use CIFAR-10 (krizhevsky2009learning) as the in-distribution data ($\mathbb{P}_{\text{in}}$). For the covariate-shifted data ($\mathbb{P}_{\text{out}}^{\text{covariate}}$), we use CIFAR-10-C (hendrycks2018benchmarking) with Gaussian additive noise for our main experiments, and provide ablations in Appendix F on other types of covariate shifts. For semantic-shifted OOD data ($\mathbb{P}_{\text{out}}^{\text{semantic}}$), we use natural image datasets: SVHN (netzer2011reading), Textures (cimpoi2014describing), Places365 (zhou2017places), LSUN-Crop (yu2015lsun), and LSUN-Resize (yu2015lsun). Large-scale results on the ImageNet dataset can be found in Section 4.4. Additional results on the PACS dataset (li2017deeper) from DomainBed is presented in Appendix E. We provide a detailed description of the datasets in Appendix C.
To simulate the wild data $\mathbb{P}_{\text{wild}}$, we mix a subset of ID data ($\mathbb{P}_{\text{in}}$) with the covariate shifted dataset ($\mathbb{P}_{\text{out}}^{\text{covariate}}$) under various $\pi_{c}\in\{0.0,0.1,0.2,0.5,0.9\}$; we default to $\pi_{c}=0.5$ for the main experiment. We keep $\pi_{s}=0.1$ to reflect the fact that we expect semantic OOD shifts ($\mathbb{P}_{\text{out}}^{\text{semantic}}$) to be encountered less frequently, which is the same value as used in (katz2022training). We split ID datasets into two halves: we use 50% as ID training data and 50% for creating the mixture data.
@@
Training details. For CIFAR experiments and methods, we use the Wide ResNet (zagoruyko2016wide) architecture with 40 layers and widen factor of 2. The model is optimized using stochastic gradient descent with Nesterov momentum (duchi2011adaptive). We set the weight decay as 0.0005, and momentum as 0.09. We initialize the model with a pre-trained network on CIFAR-10, and then trained for 100 epochs using our method. The initial learning rate is 0.0001 and decays by a factor of 2 at epochs 50, 75, and 90. Following the previous literature (katz2022training), we set $\alpha=0.05$, and set $\tau$ to be twice the loss of the pre-trained model. For all the experiments, we use a batch size of 128 and a dropout rate of 0.3. Our framework was implemented with PyTorch 1.8.1. All training is performed using NVIDIA GeForce RTX 2080 Ti. See Appendix B for additional experimental details, including our validation strategy for selecting $\eta$.
@@ 4.2 Results and Discussion
[…]
@@
Table 3: Main results: comparison with competitive OOD generalization and OOD detection methods on CIFAR-10. We run our method 3 times and report the average and std. For experiments using $\mathbb{P}_{\mathrm{wild}}$, we set $\pi_{s}=0.5$, $\pi_{c}=0.1$. For each semantic OOD dataset, we create corresponding wild mixture distribution $\mathbb{P}_{\text{wild}}:=(1-\pi_{s}-\pi_{c})\mathbb{P}_{\text{in}}+\pi_{s}\mathbb{P}_{\text{out}}^{\text{semantic}}+\pi_{c}\mathbb{P}_{\text{out}}^{\text{covariate}}$ for training and evaluating on the corresponding test dataset. $\pm x$ denotes the standard error, rounded to the first decimal point. Results for LSUN-R and Texture datasets are in Appendix D. (*Since all the OOD detection methods use the same model trained with the CE loss on $\mathbb{P}_{\text{in}}$, they display the same ID and OOD accuracy on CIFAR-10-C.)
@@
[…]
@@
We highlight a few observations: (1) Scone outperforms competitive post hoc OOD detection methods, including MSP (hendrycks2016baseline), ODIN (liang2018enhancing), Energy (liu2020energy), Mahalanobis (lee2018simple),  ViM (wang2022vim), and the latest baseline KNN (sun2022out) — all of which use the same model trained with the CE loss on $\mathbb{P}_{\text{in}}$ (and hence they display the same OOD accuracy on CIFAR-10-C). As an example of our method’s improved performance, when using SVHN as $\mathbb{P}_{\text{out}}^{\text{semantic}}$ our method yields an FPR95 of 10.86% (lower is better), which outperforms the best baseline performance of 12.89%. At the same time, the OOD generalization performance is significantly improved from 75.05% (baseline CE model) to 84.69% (ours). (2) Our method also outperforms common OOD generalization baselines, including IRM (arjovsky2019invariant), Mixup (zhang2018mixup), and VREx (krueger2021out). While these methods display stronger OOD generalization performance than the ERM (or CE) baseline, they underperform ours both in terms of OOD generalization and OOD detection. [^2] Same as ours, we use the energy score in test time to perform OOD detection. (3) Lastly, we also compare to OOD detection methods utilizing $\mathbb{P}_{\text{wild}}$, including oe (hendrycks2018deep), energy-regularized learning (liu2020energy), and WOODS (katz2022training), which is the latest such method to be developed. These methods are among the strongest OOD detection methods, yet display a significantly worsened OOD generalization performance. The main reason is that they make assumptions on $\mathbb{P}_{\text{wild}}$ without considering covariate OOD data.
@@
[…]
@@
In Table 4, we ablate the effect of $\pi_{c}$, which modulates the fraction of covariate OOD data in the mixture distribution $\mathbb{P}_{\text{wild}}$. For all settings, we contrast the performance without vs. with margin. The margin in each $\pi_{c}$ setting is validated using the strategy in Appendix B. Here we consistently use $\pi_{s}=0.1$, which reflects the practical scenario that the majority of test data may remain in the known classes. We primarily focus on evaluations where $\pi_{c}\neq 0$, since our problem setting uniquely introduces the covariate shift in the wild distribution. However, for completeness, we also include results when $\pi_{c}=0$. We highlight a few interesting observations: (1) without any enforced margin, the OOD generalization performance for WOODS (katz2022training) generally degrades with increasing $\pi_{c}$. This is likely due to the fact that a larger $\pi_{c}$ translates into more severe covariate shifts. For example, when $\pi_{c}=0.9$, the OOD classification accuracy decreases to 52.84%. (2) Our method is overall more robust under large $\pi_{c}$ settings than the WOODS baseline. For instance, in a challenging case with $\pi_{c}=0.9$, Scone outperforms WOODS by 33.34%. Overall, these results demonstrate the benefits of Scone for both OOD generalization and detection.
@@
WOODS (katz2022training) 44.46 86.49 10.50 98.22